# Trajectory Optimization Under Stochastic Dynamics Leveraging Maximum Mean Discrepancy

### v1

@@ I Introduction
Risk-aware trajectory optimization provides a rigorous template for assuring safety under stochastic dynamics model [1], [2]. In this paper, we adopt the premise of sampling-based optimization that relies on simulating the stochastic dynamics [1], [3]. These class of approaches can be applied to arbitrary noise model and even non-differentiable black-box functions. As a result, they allow us to relax the linear dynamics and Gaussian noise assumption prevalent in many existing works, e.g, [4], [5].
@@
In this paper, we consider a setting, where cost/constraint evaluations along the state trajectory samples are computationally expensive, e.g due to expensive collision checks. Furthermore, often, computing the rollouts itself could be computationally and memory intensive. Our key idea is to distill the statistical information contained in $\tilde{N}$ state trajectory samples into a reduced-set with sample size $N<<\tilde{N}$. For example, in our implementation, $\tilde{N}\approx N^{2}$. Furthermore, the distillation to reduced-set should be done in a way that it ensures reliable downstream risk estimation using only the $N$ samples.
Algorithmic Contribution: We formalize our key ideas using distribution embedding in Reproducing Kernel Hilbert Space (RKHS) and the associated Maximum Mean Discrepancy (MMD) measure [6],[7]. Specifically, we formulate a MMD-based surrogate for state dependent risk. The important feature of our risk surrogate is that we can apply the tools from RKHS distribution embedding to improve its sample efficiency. Put differently, we propose a bi-level optimization that provides a systematic way of reducing the number of state trajectory samples needed to estimate the risk associated with a sequence of control commands. Finally, we present a custom sampling-based optimizer to minimize the proposed risk surrogate.
State-of-the-Art Performance: We perform two sets of benchmarking to demonstrate the advantages of our approach. First, we show that our MMD-based surrogate is more sample efficient in estimating collision risk in a given scene, as compared to popular alternatives based on CVaR [1],[8],[9],[10],[11]. The difference is particularly stark in the low sample regime. Second, we also apply our proposed trajectory optimizer in a Model Predictive Control(MPC) setting in a high-fidelity simulator CARLA[12] to highlight the improvements over a deterministic noise-ignorant baseline as well as a CVaR-based baseline.
@@ II-A Trajectory Optimization
where $\mathbf{x}_{k}=(s_{k},d_{k},\psi_{k},\dot{\psi}_{k},v_{k})$ represents the state of the vehicle at time-step $k$. The vector $\mathbf{x}$ is the concatenation of the states at different $k$. The function $f$ is taken from [13]. The variable $v_{k}$ is the longitudinal velocity of the ego-vehicle. The control inputs are the longitudinal acceleration $a_{k}$ and steering input $\theta_{k}$. The vectors $\mathbf{a}$ and $\boldsymbol{\theta}$ are formed by stacking $a_{k}$ and $\theta_{k}$ at different time-step $k$ respectively. The stochastic disturbances acting on the vehicle are modeled as an effect of the perturbation of the nominal acceleration and steering inputs by $\epsilon_{a,k}$ and $\epsilon_{\theta,k}$ respectively. Let $\boldsymbol{\epsilon}_{a}$, $\boldsymbol{\epsilon}_{\theta}$ be vectors formed by stacking $\epsilon_{a,k}$ and $\epsilon_{\theta,k}$ respectively at different $k$. We assume that $\boldsymbol{\epsilon}_{a}\sim p_{a}$, $\boldsymbol{\epsilon}_{\theta}\sim p_{\theta}$. The perturbation at the control level is mapped to the state trajectory distribution $p_{\mathbf{x}}$ through $f$. For the sake of generality, we assume that $p_{a}$ and $p_{\theta}$ are dependent on the control inputs. For example, the slippage of a vehicle on icy-roads depends on the magnitude of the acceleration and steering commands. We don’t make any assumptions on the parametric form of $p_{a},p_{\theta}$ and $p_{\mathbf{x}}$. Instead, we just rely on the ability to sample from $p_{a},p_{\theta}$ and rollout the dynamics for every sampled ${\epsilon}_{a,k},{\epsilon}_{\theta,k}$.
The first term in the cost function (1a) minimizes the expected cost over the state distribution. Typically, these would contain costs such as path-following error for which average performance is satisfactory. In contrast, the second term captures the risk associated with state trajectory distribution $\mathbf{x}$ and depends on its higher-order noise characteristics like variance, skewness, kurtosis, etc. Intuitively, the risk captures the probability of a occurrence of an event (e.g collision) for a given sequence of control inputs. We present the algebraic form our risk model in the next subsection. The last terms in (1a) penalizes the use of large control inputs. Each term in the cost function is weighted by $w_{i}$ that can be used to tune the risk-seeking behavior of the ego-vehicle. We also consider control bounds through inequalities (1c).
@@ III-A State Risk as Difference of Distributions
In the deterministic scenario, driving $\overline{h}(\mathbf{x})$ to zero will push $h(\mathbf{x})$ to the feasible boundary. In the stochastic case, (3) maps $p_{\mathbf{x}}$ to a distribution of constraint residuals. Let $\overline{h}(\mathbf{x})\sim p_{\overline{h}}$. The key insight in our work is that although we don’t know the parameteric form for $p_{\overline{h}}$, we can be certain that its entire mass lies to the right of $\overline{h}=0$ (see Fig.2). Moreover, as $P(h(\mathbf{x})\geq 0)$ approaches zero, the $p_{\overline{h}}$ converges to a Dirac-Delta distribution $p_{\delta}$. In other words, one way of reducing risk is to minimize the difference between $p_{\overline{h}}$ and $p_{\delta}$. Thus, we propose the following risk estimate following our prior works [15], [16].
@@
where $\mathcal{L}_{dist}$ is any measure that characterizes the difference between two distributions. In the subsequent sections, we propose MMD as a potential choice for $\mathcal{L}_{dist}$.
@@ III-C MMD Based Risk Surrogate
Why is $r_{MMD}^{emp}$ an Ideal Choice for Risk Estimate? Eqn.(7b) suggests that we need $N^{2}$ rollouts of state trajectory $\mathbf{x}$ and evaluate $\overline{h}(\mathbf{x})$ over each of them to compute $\hat{\mu}[\overline{h}]$ and consequently $r_{MMD}^{emp}$. However, this can be computationally prohibitive if evaluating $\overline{h}$ is difficult, e.g, due to expensive collision checks. Fortunately, RKHS embedding and MMD provide us a set of tools to systematically choose only $N$ out of those $N^{2}$ rollouts to estimate $\overline{h}(\mathbf{x})$ and $r_{MMD}^{emp}$ and yet ensure minimal loss in risk estimation accuracy.
@@
Now, Theorem 1 from [6] ensures that if $\hat{\mu}[\mathbf{x}^{^{\prime}}]$ is close to $\hat{\mu}[\mathbf{x}]$ in MMD sense, then $\hat{\mu}[\overline{h}^{^{\prime}}]$ will be close to $\hat{\mu}[\overline{h}]$ in the same metric. Consequently, $r_{MMD}^{emp}$ computed on the smaller $N$ rollouts will be close to that computed over the larger $N^{2}$ samples. Alternately, if we can minimize $\left\|\hat{\mu}[\mathbf{x}^{^{\prime}}]-\hat{\mu}[\mathbf{x}]\right\|_{\mathcal{H}}^{2}$ , then we can ensure minimal loss in risk estimation accuracy while reducing the sample-size. This is discussed in the next subsection.
@@ III-E Trajectory Optimizer
Alg.1 presents our approach for solving (1a)-(1c) when the risk cost is given by $r_{MMD}^{emp}$. Our approach combines concepts from constrained version of gradient-free Cross Entropy Method (CEM) [18] and, Model Predictive Path Integral (MPPI) [20] and convex optimization. The basic idea is to use sampling to iteratively arrive at control inputs that have low risk and expected cost as well as small magnitudes.
Alg.1 begins by initializing the mean and covariance for sampling in Line 2. However, instead of directly sampling control inputs, we sample longitudinal velocity and lateral-offsets (from reference path) set-points from a Gaussian distribution (Line 5). This is then fed to a Frenet planner (Line 6) inspired from [21] that generates a batch of trajectories. These trajectories are then converted to acceleration and control inputs by using the differential flatness property of a car-like vehicle [22]. We also clip the magnitudes to satisfy the control bounds. In Line 7, we sample the control perturbation and perform $N^{2}$ rollouts of discrete-time dynamics (1b) to obtain as many state trajectory samples for every control input sample. Note that this rollout process is repeated for every sample in the batch of control input generated in line 6. In line 8, we invoke the optimal reduced-set optimization to choose $N$ subset out of the $N^{2}$ state trajectory samples. In line 9, we estimate $r_{MMD}^{emp}$ over the reduced-set samples. In line 10, we form a $ConstraintElliteSet$ that stores the $n_{c}$ control inputs and Frenet parameters from the total batch that resulted in lowest $r_{MMD}^{emp}$. In line 11, we evaluate the risk, expected and control cost over the control samples and state trajectories stored in $ConstraintElliteSet$. In line 12, we append the costs computed over $ConstraintElliteSet$ to a list. In line 14, we form $ElliteSet$ by choosing $n_{e}$ samples of control and state trajectory from $ConstraintElliteSet$. In line 15, we update the sampling distribution of Frenet parameters through (13a)-(13b), using the samples from the $ElliteSet$.
@@
In (13c), $\gamma$ is the so-called temperature parameter of MPPI [20]. The constant $\eta$ is the learning-rate to prevent sharp changes in the mean and covariance of the sampling distribution. One advantage of Alg.1 is that we have parametrized control sampling over a long horizon in terms of low dimensional vectors of velocity and lateral offset-setpoints. This in turn improves the computational efficiency of Alg.1
@@ 19       $cost\leftarrow$ $c_{q}$, $\forall q$ in the $ConstraintEliteSet$
IV Connections to Existing Works
@@
Linear Dynamics and Constraints Under Gaussian Noise: For linear dynamics perturbed by Gaussian noise and affine per-step constraint function $h_{k}$, the risk as defined in (2) has an exact convex reformulation [5]. This ensures that the resulting trajectory optimization can be efficiently solved. For non-linear systems and constraint functions, one can resort to linearization to achieve similar structure [4], [23]. However, the state and constraint distribution resulting from linearized system may be inaccurate and compromise the resulting safety guarantees. Recently, authors in [24] presented tractable reformulations for $r$ under Gaussian Mixture Model. Our work differs from all these cited works in the sense that it puts no assumptions on the uncertainty, constraints or dynamics. In fact, our approach can work when all these system specifications are black-box.
@@
Arbitrary Dynamics, Constraints and Noise Model: Our proposed work falls in this category. Typically, when dealing with non-linear dynamics and constraint functions in their exact form (without linearization), under arbitrary noise, the risk $r$ does not have an analytical form. Thus, existing works focus on approximating risk through samples drawn from the state distribution and evaluating the constraint function on them. For example, [8], [9],[10],[11] approximate $r$ through Conditional Value at Risk (CVaR) estimate over the samples of $h({{}^{ij}}\mathbf{x})$. Authors in [1] follow a similar approach but define $r$ as CVaR over the constraint residual function $\overline{h}({{}^{ij}}\mathbf{x})$. Another possibility is to represent $r$ through the so-called sample average approximation (SAA) of $\overline{h}({{}^{ij}}\mathbf{x})$ samples [14].
@@
A key differentiating factor between the existing CVaR [1], [8], [9],[10],[11], SAA [14] based approximations and our MMD-based risk surrogate, is that we can maximize the expressive capacity of the latter for a given sample size (recall (11a)-(LABEL:bi_level_inner_red_cost)). We are not aware of any such mechanism being presented in existing works for CVaR and SAA based risk approximations. As shown in the next section, this directly translates to lower constraint violations even while using very few number of rollouts and even fewer constraint evaluations.
@@
Connections to Works on RKHS Embedding: Our formulation builds upon the RKHS embedding approach presented in [6], [17] for functions of multiple random variables. These works also suggest using the concept of reduced-set to maximize the expressive capacity of the embeddings. However [6], [17] relies on randomly selecting a subset of the samples as the reduced-set. In contrast, our bi-level optimization (11a)-(LABEL:bi_level_inner_red_cost) provides a much more principled approach. Our proposed work also extends the MMD-based risk surrogate presented in [15], [16] to the case of stochastic dynamics.

### v2

@@ I Introduction
Risk-aware trajectory optimization provides a rigorous template for assuring safety under stochastic dynamics model [1], [2]. There are two key challenges in this regard. First, characterizing the state transition distribution for non-linear systems under arbitrary noise model is often intractable. This in turn, also prevents obtaining analytical optimizer friendly expression for the underlying safety (collision, lane violation) risk. Some existing works linearize the non-linear dynamics and adopt Gaussian noise model to by-pass this intractability [3], [4]. However, such approximations can lead to incorrect state transition distribution and consequently poor collision risk estimate. Thus, in this paper, we adopt the premise of sampling-based optimization that relies on simulating the stochastic dynamics [1], [5]. These class of approaches can be applied to arbitrary noise model and even non-differentiable black-box constraint functions. As a result, they allow us to relax the linear dynamics and Gaussian noise assumption prevalent in many existing works, e.g, [3], [6].
@@
In this paper, we consider a setting, where cost/constraint evaluations along the state trajectory samples are computationally expensive, e.g due to expensive collision checks. Furthermore, often, computing the rollouts itself could be computationally and memory intensive. Our key idea is to distill the statistical information contained in $\tilde{N}$ state trajectory samples into a reduced-set with sample size $N<<\tilde{N}$. For example, in our implementation, $\tilde{N}\approx N^{2}$. The costly constraint evaluations are then performed only on the reduced-set samples. The main challenge is that the distillation to reduced-set should be done in a way that it ensures reliable downstream risk estimation using only the $N$ samples.
Algorithmic Contribution: We use distribution embedding in Reproducing Kernel Hilbert Space (RKHS) and the Maximum Mean Discrepancy (MMD) measure [7],[8] to formalize our key ideas. Specifically, we develop an MMD-based surrogate for state-dependent risk, leveraging RKHS tools to enhance sample efficiency. Our approach involves a bi-level optimization that systematically reduces the number of state trajectory samples required to estimate the risk of a sequence of control commands. Additionally, we introduce a custom sampling-based optimizer to minimize the proposed risk surrogate.
State-of-the-Art Performance: We conduct two benchmarking sets to demonstrate the advantages of our approach. First, we show that our MMD-based surrogate is more sample-efficient in estimating collision risk in a given scene compared to popular CVaR-based alternatives [1],[9],[10],[11],[12], especially in the low sample regime. Second, we apply our trajectory optimizer in a Model Predictive Control (MPC) setting within the high-fidelity CARLA simulator [13], highlighting improvements over both a deterministic noise-ignorant baseline and a CVaR-based baseline.
@@ II-A Trajectory Optimization
where $c(.)$ represents the state-dependent cost function. The vector $\mathbf{x}_{k}=(s_{k},d_{k},\psi_{k},\dot{\psi}_{k},v_{k})$ represents the state of the vehicle at time-step $k$. The vector $\mathbf{x}$ is the concatenation of the states at different $k$. The function $f$ is taken from [14]. The variable $v_{k}$ is the longitudinal velocity of the ego-vehicle. The control inputs are the longitudinal acceleration $a_{k}$ and steering input $\theta_{k}$. The vectors $\mathbf{a}$ and $\boldsymbol{\theta}$ are formed by stacking $a_{k}$ and $\theta_{k}$ at different time-step $k$ respectively. The stochastic disturbances acting on the vehicle are modeled as an effect of the perturbation of the nominal acceleration and steering inputs by $\epsilon_{a,k}$ and $\epsilon_{\theta,k}$ respectively. Let $\boldsymbol{\epsilon}_{a}$, $\boldsymbol{\epsilon}_{\theta}$ be vectors formed by stacking $\epsilon_{a,k}$ and $\epsilon_{\theta,k}$ respectively at different $k$. We assume that $\boldsymbol{\epsilon}_{a}\sim p_{a}$, $\boldsymbol{\epsilon}_{\theta}\sim p_{\theta}$. The perturbation at the control level is mapped to the state trajectory distribution $p_{\mathbf{x}}$ through $f$. For the sake of generality, we assume that $p_{a}$ and $p_{\theta}$ are dependent on the control inputs. For example, the slippage of a vehicle on icy-roads depends on the magnitude of the acceleration and steering commands. We don’t make any assumptions on the parametric form of $p_{a},p_{\theta}$ and $p_{\mathbf{x}}$. Instead, we just rely on the ability to sample from $p_{a},p_{\theta}$ and rollout the dynamics for every sampled ${\epsilon}_{a,k},{\epsilon}_{\theta,k}$.
The first term in (1a) minimizes the expected state cost, typically addressing path-following errors where average performance suffices. The second term captures risk in the state trajectory $\mathbf{x}$, considering higher-order noise characteristics like variance, skewness, and kurtosis. Intuitively, it represents the probability of an event (e.g., collision) for a given control sequence. The last term in (1a) penalizes large control inputs, with weights $w_{i}$ tuning the ego-vehicle’s risk-seeking behavior. Control bounds are enforced through (1c).
@@ III-A State Risk as Difference of Distributions
In the deterministic scenario, driving $\overline{h}(\mathbf{x})$ to zero will push $h(\mathbf{x})$ to the feasible boundary. In the stochastic case, (3) maps $p_{\mathbf{x}}$ to a distribution of constraint residuals. Let $\overline{h}(\mathbf{x})\sim p_{\overline{h}}$. The key insight in our work is that although we don’t know the parametric form for $p_{\overline{h}}$, we can be certain that its entire mass lies to the right of $\overline{h}=0$ (see Fig.2). Moreover, as $P(h(\mathbf{x})\geq 0)$ approaches zero, the $p_{\overline{h}}$ converges to a Dirac-Delta distribution $p_{\delta}$. In other words, one way of reducing risk is to minimize the difference between $p_{\overline{h}}$ and $p_{\delta}$. Intuitively this minimization will make $p_{\overline{h}}$ look similar to $p_{\delta}$. Thus, we propose the following risk estimate following our prior works [16], [17].
@@
where $\mathcal{L}_{dist}$ is any measure that characterizes the difference between two distributions. For example, Kullback-Leibler Divergence (KLD) quantifies distribution similarity but requires known analytical forms, making it unsuitable for comparing $p_{\overline{h}}$ and $p_{\delta}$ using only sample-level information. In the following sections, we propose MMD as a potential choice for $\mathcal{L}_{dist}$.
@@ III-C MMD Based Risk Surrogate
Advantages of $r_{MMD}^{emp}$ as a Risk Estimate: Eqn.(7b) suggests that we need $N^{2}$ rollouts of state trajectory $\mathbf{x}$ and evaluate $\overline{h}(\mathbf{x})$ over each of them to compute $\hat{\mu}[\overline{h}]$ and consequently $r_{MMD}^{emp}$. However, this can be computationally prohibitive if evaluating $\overline{h}$ is difficult, e.g, due to expensive collision checks. Fortunately, RKHS embedding and MMD provide us a set of tools to systematically choose only $N$ out of those $N^{2}$ rollouts to estimate $\overline{h}(\mathbf{x})$ and $r_{MMD}^{emp}$ and yet ensure minimal loss in risk estimation accuracy.
@@
Now, Theorem 1 from [7] ensures that if $\hat{\mu}[\mathbf{x}^{^{\prime}}]$ is close to $\hat{\mu}[\mathbf{x}]$ in MMD sense, then $\hat{\mu}[\overline{h}^{^{\prime}}]$ will be close to $\hat{\mu}[\overline{h}]$ in the same metric. Consequently, $r_{MMD}^{emp}$ computed on the smaller $N$ rollouts will be close to that computed over the larger $N^{2}$ samples. Alternately, if we can minimize $\left\|\hat{\mu}[\mathbf{x}^{^{\prime}}]-\hat{\mu}[\mathbf{x}]\right\|_{\mathcal{H}}^{2}$ , then we can ensure minimal loss in risk estimation accuracy while reducing the sample-size. It is worth pointing out that such sample optimization feature is not available in risk metrics like CVaR.
@@ III-E Trajectory Optimizer
Alg.1 presents our approach for solving (1a)-(1c) when the risk cost is given by $r_{MMD}^{emp}$. It combines constrained gradient-free Cross Entropy Method (CEM) [19], Model Predictive Path Integral (MPPI) [21], and convex optimization to iteratively refine low-risk, low-cost control inputs.
The algorithm initializes the sampling distribution (Line 2) and samples longitudinal velocity and lateral offset setpoints (Line 5), which are fed to a Frenet planner [22] (Line 6) to generate trajectories. Using differential flatness [23], these are converted to accelerations and clipped to control bounds. Control perturbations are sampled, and $N^{2}$ rollouts of (1b) yield state trajectory samples (Line 7). A reduced-set of $N$ trajectories is selected (Line 8), and $r_{MMD}^{emp}$ is estimated (Line 9). The lowest-risk $n_{c}$ samples form $ConstraintElliteSet$ (Line 10), from which costs are computed (Line 11) and stored (Line 12). A final $ElliteSet$ of $n_{e}$ samples is selected (Line 14) to update the sampling distribution (Line 15) via:
@@
Here, $\gamma$ is the MPPI temperature parameter [21], and $\eta$ controls learning rate. By parameterizing long-horizon control sampling using low-dimensional velocity and offset setpoints, Alg.1 significantly improves computational efficiency.
@@ 19       $cost\leftarrow$ $c_{q}$, $\forall q$ in the $ConstraintEliteSet$
III-F Diagonal Estimation of RKHS Embedding
The RKHS embedding computed in (6) is defined in the product space formed with $N$ samples each of noise $\boldsymbol{\epsilon}_{a}$ and $\boldsymbol{\epsilon}_{\theta}$. For example, if $(^{1}\boldsymbol{\epsilon}_{a},{{}^{2}}\boldsymbol{\epsilon}_{a})$ and $(^{1}\boldsymbol{\epsilon}_{\theta},{{}^{2}}\boldsymbol{\epsilon}_{\theta})$ are respectively samples of $\boldsymbol{\epsilon}_{a}$ and $\boldsymbol{\epsilon}_{\theta}$, then we form noise pairs such as $(^{1}\boldsymbol{\epsilon}_{a},^{1}\boldsymbol{\epsilon}_{\theta})$, $(^{2}\boldsymbol{\epsilon}_{a},^{1}\boldsymbol{\epsilon}_{\theta})$, etc. This in turn results in $N^{2}$ state trajectory rollouts. However, it is possible to use the so-called diagonal estimation of RKHS embedding that requires only $N$ state trajectory rollouts [7]. It is given by:
$\hat{\mu}_{D}[\mathbf{x}]=\sum_{i=1}^{i=N}\frac{1}{N}\phi({{}^{i}}\mathbf{x}),\qquad{{}^{i}}\mathbf{x}=\mathbf{x}({{}^{i}}\boldsymbol{\epsilon}_{a},{{}^{i}}\boldsymbol{\epsilon}_{\theta})$ (14)
The embedding $\hat{\mu}_{D}[\mathbf{x}]$ (14) has a higher variance than $\hat{\mu}[\mathbf{x}]$ that also translates to the MMD-based risk estimation. We will refer to the risk cost computed using $\hat{\mu}_{D}[\mathbf{x}]$ as $r_{MMD-D}^{emp}$. Nevertheless, in very low sample regimes, $\hat{\mu}_{D}[\mathbf{x}]$ can often give competitive performance and we analyze this further in Section V.
IV Connections to Related Works
@@
Linear Dynamics and Constraints Under Gaussian Noise: For linear dynamics perturbed by Gaussian noise and affine per-step constraint function $h_{k}$, the risk defined in (2) has an exact convex reformulation [6], allowing for efficient trajectory optimization. For non-linear systems and constraint functions, linearization can achieve a similar structure [3], [24], but this may lead to inaccuracies in the state and constraint distribution, compromising safety guarantees. Recently, [25] introduced tractable reformulations for $r$ under Gaussian Mixture Models. In contrast, our work only need sample level information of the uncertainty and imposes no assumptions on the algebraic form of the constraints, or dynamics.
@@
Arbitrary Dynamics, Constraints and Noise Model: In this context, the risk $r$ lacks an analytical form, leading existing works to approximate it using samples drawn from the state distribution to evaluate the constraint function. For instance, [9], [10], [11], and [12] estimate $r$ through Conditional Value at Risk (CVaR) over samples of $h({{}^{ij}}\mathbf{x})$. Similarly, [1] defines $r$ as CVaR over the constraint residual function $\overline{h}({{}^{ij}}\mathbf{x})$. Another approach involves representing $r$ through the sample average approximation (SAA) of $\overline{h}({{}^{ij}}\mathbf{x})$ samples [15].
@@
A key differentiating factor between existing CVaR [1], [9], [10], [11], [12], and SAA [15] approximations and our MMD-based risk surrogate is our ability to maximize the expressive capacity of the latter for a given sample size (see (11a)-(LABEL:bi_level_inner_red_cost)). We are not aware of any similar mechanism in existing works for CVaR and SAA-based risk approximations.
Chance-Constrained Optimization: Instead of minimizing risk $r$, we can enforce constraints of the form $r\leq\gamma$ for some $\gamma\in[0,1]$, leading to a more restrictive chance-constrained optimization setting. These problems are typically tractable only for linear dynamics, Gaussian uncertainty, and affine $h(\mathbf{x})$, while more general cases often rely on CVaR-based reformulations [9]. As discussed in Section III, our MMD-based risk serves as a surrogate for $r$, aiming to minimize it rather than enforcing an upper bound, which may be difficult to determine beforehand.
@@
Connections to Works on RKHS Embedding: Our formulation builds upon the RKHS embedding approach presented in [7], [18] for functions of multiple random variables. These works also suggest using the concept of reduced-set to maximize the expressive capacity of the embeddings. However [7], [18] relies on randomly selecting a subset of the samples as the reduced-set. In contrast, our bi-level optimization (11a)-(LABEL:bi_level_inner_red_cost) provides a much more principled approach. Our proposed work also extends the MMD-based risk surrogate presented in [16], [17] to the case of stochastic dynamics. A slightly different perspective from our approach is the setting of Distributionally Robust Optimization (DRO). It is typically employed when the underlying probability distribution is itself not known accurately and operates by by optimizing over an ambiguity set of possible distributions. In contrast, our work assumes a well-estimated noise distribution for direct risk modeling. However, our method can be extended to DRO using MMD-based ambiguity sets [26], offering a promising avenue for future research.